OpenStudy (anonymous):

find all real numbers roots of each equation log base 3 [log base3 (2x)]=-2 I can't remember how to do these can someone show me?

jhonyy9 (jhonyy9):

- so there is log3 (2x) =-2 ???

jhonyy9 (jhonyy9):

- if is like so than will be : log3 (2x)=-2 log3 (2x)=-2log3 (3) log3 (2x)=log3 (3)^(-2) 2x=3^(-2) 2x=1/3^2 2x=1/9 x=1/9/2 x=1/9 * 1/2 x=1/18